A solution is 0.100 M in substances A and B, which react by a first-order reaction. If the reaction is 15.0%
complete in 6.75 min, what is its half-life under these conditions? What would be the half-life if A and B
were each at 0.200 M, and how long would it take for 15.0% completion of the reaction?
1
Expert's answer
2020-04-09T13:11:09-0400

v=kC - rate of a first-order reaction. k-rate constant of reaction.

v=-dC/dt

-dC/dt=kC

dC/C=-kdt

"\int"dC/C=-"\int"kdt

lnC-lnC0=-kt оr ln(C/C0) =-kt

C/C0=0.85 (C=C0-0.15C0=0.85C0); t=6.75min (the reaction is 15.0% complete in 6.75 min)

ln(0.85)=-k*6.75min

k=-ln(0.85) / 6.75min=0.024min-1

t=-ln(C/C0) / k

T1/2=-ln((C0/2)/C0)/k or T1/2=ln2/k - the half-life does not depend on initial concentration (as well as the 15% completion time of the reaction, as well as the time of any degree of completion of first-order reactions)

T1/2=0.693/k=0.693/0.024min-1=28,88min.

For the above reasons, the half-life of A and B at an initial concentration of 0.200M, and the reaction completion time of 15.0% will be the same as for a concentration of 0.100M.